OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (anonymous):

so your slope in the equation is -1. However, a perpendicular line will have a negative reciprocal of the slope. So the negative reciprocal of -1 is 1. now lets put your info into y = mx + b using slope 1 and points (2,-2). y = mx + b -2 = 1(2) + b -2 = 2 + b -2 -2 = b -4 = b so your equation is y = 1x - 4 or just y = x - 4